Please help me understand this problem,The length of a rectangular court of badminton is 27 feet longer than its width. If its perimeter is 122 feet. What are the dimensions of the rectangular court? Thank you :)
Answer by ikleyn(52794)

.
Let L be the length, W be the width, in feet.


The condition says  L = W + 27.


It also says about the perimeter


L + W + L + W = 122.


Replace here L by (W + 27). You will get


(W+27) + W + (W+27) + W = 122,  or


4W + 54 = 122  ---->  4W = 122 - 54 = 68  --->  W = 68/4 = 17.


Answer.  The width of the court is 17 ft. The length is 17+ 27 = 44 ft.


Check.  44 + 17 + 44 + 17 = 122.  Correct !